2. Resolving to read more each week, Drew starts a new reading program and promises to read 10 minutes the first
week, 20 minutes the second, 30 minutes the third, and so on.
a. How much time will Drew spend reading during the 12th week?
1
Expert's answer
2020-04-09T14:09:38-0400

We can write time duration for each week like follow

First week = 10 minutes = (10 * 1) minutes

Second week = 20 minutes = (10 * 2) minutes

Third week = 30 minutes = (10 * 3) minutes

Fourth week = 40 minutes = (10 * 4) minutes

Similarly, in the nth week, he will read = 10 * n minutes

So we have to find reading time during 12th week (i.e. n = 12)

Reading time = 10 * 12 = 120 minutes

Drew will spend 120 minutes during 12th week.